# Belo Horizonte, MG, Brazil

### v3

@@ Abstract
Accurate modeling of spatial dependence is pivotal in analyzing spatial data, influencing parameter estimation and predictions. The spatial structure of the data significantly impacts valid statistical inference. Existing models for areal data often rely on adjacency matrices, struggling to differentiate between polygons of varying sizes and shapes. Conversely, data fusion models rely on computationally intensive numerical integrals, presenting challenges for moderately large datasets. In response to these issues, we propose the Hausdorff-Gaussian process (HGP), a versatile model utilizing the Hausdorff distance to capture spatial dependence in both point and areal data. Integration into generalized linear mixed-effects models enhances its applicability, particularly in addressing data fusion challenges. We validate our approach through a comprehensive simulation study and application to two real-world scenarios: one involving areal data and another demonstrating its effectiveness in data fusion. The results suggest that the HGP is competitive with specialized models regarding goodness-of-fit and prediction performances. In summary, the HGP offers a flexible and robust solution for modeling spatial data of various types and shapes, with potential applications spanning fields such as public health and climate science.
Keywords: CAR models; Geostatistics; Spatial data fusion; Spatial modeling; Spatial prediction
@@ 1 Introduction
Accounting for spatial dependence is pivotal in spatial data analysis. Valid statistical inference are deeply influenced by the spatial structure and geometry of the data. Point-referenced data are typically analyzed using geostatistics methods (e.g., Cressie,, 1993). For areal data, the conditional autoregressive model (Besag,, 1974, CAR) and its variants (Besag et al.,, 1991; Rodrigues and Assunçao,, 2012; Datta et al.,, 2019; Cruz-Reyes et al.,, 2023) are often preferred. The nature of the observed spatial data significantly influences the choice of statistical methodology. This becomes increasingly critical in complex inferential problems, such as predicting a variable in a target map using data from a source map. This issue, known as change of support (Cressie,, 1993, pg. 284–289), requires tailored methods depending on the spatial units involved (Gelfand et al.,, 2001). Spatial data from diverse sources often differ in spatial units or supports, resulting in spatial misalignment (Gotway and Young,, 2002). Data fusion is necessitated when a spatial variable is provided by multiple, potentially misaligned, sources (Moraga et al.,, 2017).
@@
Both traditional areal models and data fusion models for spatial data exhibit inherent limitations. Traditional areal models face challenges with predictions for overlapping spatial units (Bakar and Jin,, 2020), do not account for region size or shape of the areal units and lack interpretable spatial dependence parameters (Assunção and Krainski,, 2009; Cruz-Reyes et al.,, 2023). An exception of the last limitation is the directed acyclic graph auto-regressive (DAGAR) model, which offers an interpretable parameter (Datta et al.,, 2019). Data fusion models, which view areal data as averages of a continuous process, may be unsuitable for discrete outcomes (Gelfand et al.,, 2001). Additionally, their reliance on arbitrary discretization for likelihood evaluation lacks consensus on optimal resolution (Fuentes and Raftery,, 2005; Liu et al.,, 2011; Wang and Furrer,, 2021) and can introduce unquantifiable biases (Gonçalves and Gamerman,, 2018).
@@
We introduce a unified and flexible approach for analyzing spatial data of varying types, encompassing areal, point-referenced, and mixed spatial supports. Our key objective is to develop a model capable of seamlessly handling these diverse data formats within a single framework. This is achieved by combining Gaussian processes (GPs) with the Hausdorff distance (Arutyunov et al.,, 2016). The Hausdorff distance, defined on a reference metric space, effectively measures the distance between two sets while accounting for their shape. For point pairs, it reduces to the distance metric associated to the reference metric space. GPs, recognized for their flexibility, are widely utilized in spatial data modeling, notably as priors in generalized mixed effects models (Diggle et al.,, 1998), in spatially varying regression coefficients (Gelfand et al.,, 2003), and in addressing spatial misalignment issues (Banerjee and Gelfand,, 2002). We propose the Hausdorff-Gaussian process (HGP), a GP with a correlation function defined on the Hausdorff distance between spatial units. This allows us to directly incorporate shape and size information into the analysis. A crucial aspect of our work is to assess the competitiveness of the HGP, particularly when compared to specialized models designed for areal and mixed spatial data. We investigate whether the HGP, using a powered exponential correlation, can effectively model areal and mixed spatial data, thereby offering a novel and versatile tool for analyzing spatial and spatiotemporal data with mixed supports.
@@
This paper is structured as follows. Section 2 introduces the Hausdorff distance, discussing its properties, and then delineates the HGP, highlighting its significant applications. Section 3 sets up spatial generalized linear mixed models (GLMM) with HGP components and details their inferences in the Bayesian framework. A simulation study to compare the proposed model with specific models for areal and fused data is presented in Section 4. Section 5 demonstrates the practical utility of our model through two distinct applications, one on areal data and the other on data fusion. Section 6 concludes with a discussion.
@@ 2.1 Hausdorff Distance
The Hausdorff distance is a metric measuring the distance between two sets. Specifically, it is the greatest distance from a point in one set to the closest point in the other set. Consider a metric space $(D,d)$, where $D$ is a spatial region of interest and $d(x,y)$ defines the distance between any two elements $x,y\in D$. For a single point $x$ and a set $A$, define $d(x,A)=\inf_{a\in A}d(x,a)$ as the closest distance from $x$ to any element in $A$. The directed Hausdorff distance from a set $A$ to a set $B$ is
$\vec{h}(A,B)=\sup_{a\in A}d(a,B).$
This measure is not necessarily symmetric. The symmetric Hausdorff distance is defined as the greater of the two directed Hausdorff distances:
@@
Note that if $A$ and $B$ are both singletons, then $h(A,B)=d(A,B)$. Interestingly, if $A$ and $B$ are compact, there exist $a\in A$ and $b\in B$ such that $h(A,B)=d(a,b)$ (Arutyunov et al.,, 2016).
In spatial analysis, the Hausdorff distance provides a robust and suitable metric over non-empty compact subsets of $D$, denoted as $\mathcal{B}(D)$ (Molchanov,, 2005, Appendix C). While not functioning as a conventional distance in unbounded sets, it is highly effective in bounded contexts, which are prevalent in spatial statistics. Typically, the study region $D$ in spatial statistics is a bounded subset of $\mathbb{R}^{2}$, where the Hausdorff distance excels in quantifying distances between spatial units by capturing both their shapes and sizes. The metric space induced by the tuple $(\mathcal{B}(D),h)$ inherits properties like compactness and completeness from $(D,d)$. Further details regarding the topology induced by the Hausdorff distance can be found in Appendix C inMolchanov, (2005).
AB
AB
A.B
Figure 1: Three different pairs of bounded subsets of the Euclidean plane. In (a) the sets AA and $B$ share a single point on their border. In (b), they have non-empty intersection. While in (c), $B$ is a singleton set contained by $A$. The dashed lines represent the Hausdorff distance between the sets.
To demonstrate the efficacy of the Hausdorff distance in measuring distances between various spatial units, Figure 1 presents three simple examples. In each sub-figure, set $A$ is represented as a circle with a radius of 1. Figure 1a introduces a set $B$, a circle tangent to $A$. Figure 1b examines the distance between two sets having a non-empty intersection. Figure 1c features $B$ as an interior point of $A$. The dashed lines in the figures delineate the Hausdorff distances between the sets, highlighting the metric’s ability to consider the shape, size, and orientation of the objects. This contrasts with other simpler distance functions between non-empty sets, such as the infimum distance between two sets, denoted $\inf_{x\in X}\inf_{y\in Y}d(x,y)$. This distance does not fulfill the metric property of positiveness condition. In other words, $d(x,y)=0$ does not imply $x=y$. Indeed, employing such distance latter would yield a distance of zero between the three pairs of sets depicted in Figure 1.
@@ 2.2 Hausdorff–Gaussian Process
HGP is a natural extension of GP for modeling spatial data defined on the Hausdorff distance. A GP is a stochastic process over an index set such that all its finite-dimensional marginal distributions are multivariate normal distributions. An HGP expands the index set on a spatial region $D$ to encompasses a broader range of spatial units, $\mathcal{B}(D)$, which contains non-empty, bounded, and compact subsets of $D$. To define an isotropic HGP $Z(\cdot)$ over $\mathcal{B}(D)$, we specify a marginal mean function function $m(\cdot)$, a positive-valued marginal standard deviation (SD) function $v(\cdot)$, and an isotropic correlation function $r(h)$ for any two sets apart in Hausdorff distance $h$. With these functions, an HGP over $\mathcal{B}(D)$ is a stochastic process $Z(\cdot)$ whose marginal distribution at $\mathbf{s}_{1},\ldots,\mathbf{s}_{p}\in\mathcal{B}(D)$ is multivariate normal with mean vector and covariance matrix, respectively,
@@
for $i,j=1,\ldots,p$, where $h(\mathbf{s}_{i},\mathbf{s}_{j})$ is the Hausdorff distance between $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$. This process allows for representations of diverse spatial data, from areal to mixed, making it suitable for a wide range of applications such as disease mapping and data fusion.
Denote the HGP process defined in Equation (1) by $\textrm{HGP}\{m(\cdot),v(\cdot),r(\cdot)\}$. The mean function $m(\cdot)$ captures large-scale variations, usually through covariates. The SD function $v(\cdot)$ function accommodates both homoscedastic and hereroscedastic scenarios. For instance, with a log link and covariate $w(\mathbf{s})$ at spatial unit $\mathbf{s}$, a natural model is
$\log v(\mathbf{s})=\alpha_{0}+\alpha_{1}w(\mathbf{s}).$ (2)
Equation (2) allows a distinction between point-referenced and areal data, a need in data fusion applications (Wang and Furrer,, 2021). With $\mathcal{A}(\cdot)$ being the function that returns the area of a spatial object. A covariate $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$ enables different variances between these data types. Certainly, care is necessary in specifying $v(\mathbf{s})$ to ensure a valid GP (Palacios and Steel,, 2006).
@@ 2.3 Correlation Function
The correlation function $r(\cdot)$ needs to be carefully chosen to ensure the validity of the HGP as it defines the spatial dependence structure across different locations. In particular, we focus on isotropic correlation functions. Therefore, unless explicitly mentioned, all subsequent references to correlation functions will assume they are isotropic. A valid correlation function must be positive definite and meet two key criteria: $r(0)=1$ and $\lim_{h\to\infty}r(h)=0$ (Cressie,, 1993, Ch. 2.3.2), where $h$ is the Hausdorff distance between a pair $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{B}(D)$. The latter ensures that spatial correlation diminishes with increasing distance. It is noteworthy that while most geostatistical correlation functions are constructed for Euclidean spaces (Cressie,, 1993, Ch. 2.5.1), their positive definiteness may not extend to non-Euclidean spaces (Curriero,, 2006; Li et al.,, 2023).
@@
where $\nu$ controls the smoothness and $\varphi$ governs the range of the spatial dependence. To enhance interpretability, we reparametrize this function with $\rho={\log(10)}^{1/\nu}\varphi$. Consequently, $\rho$ is the distance at which the spatial correlation between a pair of spatial units drops to $0.10$, commonly referred to as the practical range. The PEC function is appealing because, for $\nu\in(0,1]$, it is positive definite as long as the distance function is conditionally negative definite (Schoenberg,, 1938). To the best of our knowledge, there exist no theoretical results on positive definite functions for the Hausdorff distance. Therefore, we recommend an empirical assessment of the PEC function positive definiteness before analyzing the data.
To this end, we empirically assess the positive definiteness of the PEC function for $\nu\in(0,1]$. We first calculate Hausdorff distance matrices for the areal and fused data to be analyzed in Sections 5.1 and 5.2, respectively. Evaluating the PEC function across these matrices with $\nu\in[0.01,1]$ and $\rho\in[d_{\min},d_{\max}]$ (where $d_{\min}$ and $d_{\max}$ are the dataset’s minimum and maximum distances) yields corresponding correlation matrices. We then compute each matrix’s smallest eigenvalue. A negative eigenvalue suggests the corresponding PEC function is not positive definite. However, even valid covariance functions can be ill-conditioned or nearly rank-deficient at dense locations or when there exists a severely strong spatial dependence (Banerjee et al.,, 2012).
@@
Figure 2 displays the smallest eigenvalues ($\lambda_{1}$) of the PEC-induced correlation matrices for both datasets. The fused data exhibits positive eigenvalues for all $(\nu,\rho)$ pairs. Conversely, the areal data reveals some small negative eigenvalues (range 0.0099-0.0099 to 0.00006-0.00006) under severely strong spatial dependence and at $\nu\approx 1$, the border of the parametric space. Therefore, our empirical study suggests the positive definiteness of the PEC function for $\nu\in(0,1]$.
@@ 3.1 Model
Generalized linear models for spatial data can benefit from replacing the GP with the HGP as spatial random effects. In this context, assume $\mathcal{S}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}$ is a set of spatial units or locations such that $\mathbf{s}_{i}\in\mathcal{B}(D)$, where $D\subset\rm I\!R^{2}$. Consider a response variable $Y(\mathbf{s}_{i})$ and a $p$-dimensional vector of covariates $\mathbf{X}_{i}$ associated to the $i$-th spatial unit. Let $\mathbf{Z}={\big[Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n})\big]}^{\top}$ be an $n$-dimensional vector of latent spatial random effects. Conditional on $Z(\mathbf{s}_{i})$ and $\mathbf{X}_{i}$, we assume that the responses $Y(\mathbf{s}_{i})$’s are mutually independent. The conditinal distribution of $Y(\mathbf{s}_{i})$ is characterized by its mean mean and possibly some additional parameters. This leads to a spatial GLMM built on an HGP.
Specifically, the spatial GLMM is specified as
$Y(\mathbf{s}_{i})\mid\mathbf{X}_{i},Z(\mathbf{s}_{i})\overset{{\rm ind}}{\sim}f(\cdot\mid\mu_{i},\bm{\gamma}),$ (3)
where $f(\cdot\mid\mu_{i},\bm{\gamma})$ is a probability density or mass function of a distribution with expectation $\mu_{i}$ and an additional parameter vector $\bm{\gamma}$. The conditional mean $\mu_{i}$ is linked to the covariate vector and the spatial random effects through a suitable, known link function $g$:
@@
where $\bm{\beta}$ is a $p$-dimensional regression coefficient vector. The spatial dependence among the responses $\mathbf{Y}={\big[Y(\mathbf{s}_{1}),\ldots,Y(\mathbf{s}_{n})\big]}^{\top}$ is introduced by an HGP on $\mathbf{Z}$, namely,
$\mathbf{Z}\sim\textrm{HGP}\big\{0,v(\cdot;\bm{\sigma}),r(\cdot;\bm{\delta})\big\},$ (5)
with zero-mean, a SD function $v(\cdot;\bm{\sigma})$ with parameters $\bm{\sigma}$, and the PEC function $r(\cdot,\bm{\delta})$ with parameters $\bm{\delta}=\{\rho,\nu\}$. To summarize, the parameters of the spatial GLMM are $\bm{\theta}={\{\bm{\beta}^{\top},\bm{\sigma}^{\top},\bm{\delta}^{\top},\bm{\gamma}^{\top}\}}^{\top}$.
A general spatial GLMM is obtained by choosing a suitable $f(\cdot\mid\mu_{i},\bm{\gamma})$ for the response. For binary response variables, a Bernoulli distribution yields a spatial logistic regression. Count data can be modeled with either a Poisson or Negative Binomial distribution, depending on the presence of overdispersion. When the conditional distribution of the response is normal, the additional parameter $\gamma$ is the measurement error or small scale variance parameter. Let $\mathbf{X}$ be a $n\times p$ matrix with its $i$-th row given by $\mathbf{X}_{i}$. From the conditional dependence, the joint density (or mass) function of $\mathbf{Y}$ conditioning on $\mathbf{X}$ and $\mathbf{Z}$ is
$p(\mathbf{y}\mid\mathbf{X},\mathbf{z})=\prod_{i=1}^{n}f(y(\mathbf{s}_{i})\mid\mu_{i},\bm{\gamma}),$ (6)
where $\mathbf{y}$ and $\mathbf{z}$ are realizations of $\mathbf{Y}$ and $\mathbf{Z}$, respectively.

### v4

# Statistical Inferences and Predictions for Areal Data and Spatial Data Fusion with Hausdorff–Gaussian Processes

@@ Abstract
Accurate modeling of spatial dependence is pivotal for both parameter estimation and prediction in spatial data analysis. The spatial structure of the data significantly narrows the set of available models. Existing areal data models often ignore polygon geometry, while data fusion models require computationally intensive integration, limiting their application. In response to these issues, we propose the Hausdorff-Gaussian process (HGP), a versatile model utilizing the Hausdorff distance to capture spatial dependence in both point and areal data. Integration into generalized linear mixed-effects models enhances its applicability, particularly in addressing data fusion challenges. We validate our approach through a comprehensive simulation study and application to two real-world scenarios: one involving areal data and another demonstrating its effectiveness in data fusion. Results show the HGP is a competitive, flexible, and robust solution for modeling diverse spatial data, with potential applications in fields like public health and climate science.
Keywords: CAR models, Geostatistics, Spatial data fusion, Spatial modeling, Spatial prediction
@@ 1 Introduction
Accounting for spatial dependence is pivotal in spatial data analysis. The nature of the observed spatial data significantly influences the choice of statistical methodology. Point-referenced data are typically analyzed using geostatistics methods (e.g., cressie1993statistics). For areal data, the conditional autoregressive model (besag1974spatial, CAR) and its variants (besag1991bayesian; rodrigues2012bayesian; datta2019spatial; cruz2023inducing) are often preferred. These differences become increasingly critical in complex inferential problems, such as predicting a variable in a target map using data from a source map. This issue, known as change of support (cressie1993statistics, pg. 284–289), requires tailored methods depending on the spatial units involved (gelfand2001change). Spatial data from diverse sources often differ in spatial units or supports, resulting in spatial misalignment (gotway2002combining). Data fusion is necessitated when a spatial variable is provided by multiple, potentially misaligned, sources (moraga2017geostatistical).
@@
Both traditional areal models and data fusion models for spatial data exhibit inherent limitations. Traditional areal models lack interpretable spatial dependence parameters (assuncao2009neighborhood). An exception of that last limitation is the directed acyclic graph auto-regressive (DAGAR) model, which offers an interpretable parameter (datta2019spatial). These models, which rely on a fixed adjacency graph, are not suited for out-of-sample prediction, as introducing new spatial units fundamentally alters the whole model structure. Furthermore, when overlapping spatial units are present (bakar2020areal), another limitation is that they do not account for the size or shape of the areal units. Data fusion models, which view areal data as averages of a continuous process, may be unsuitable for discrete outcomes (gelfand2001change). Additionally, their reliance on arbitrary discretization for likelihood evaluation lacks consensus on optimal resolution (fuentes2005model; liu2011empirical; wang2021combining) and can introduce unquantifiable biases (goncalves2018exact).
@@
We introduce a unified approach for areal, point-referenced, and mixed spatial data by combining Gaussian processes (GPs) with the Hausdorff distance (arutyunov2016some), which seamlessly handles these diverse formats in a single framework. The Hausdorff distance, defined on a reference metric space, effectively measures the distance between two sets while accounting for their shape. For point pairs, it reduces to the distance metric associated to the reference metric space. GPs, recognized for their flexibility, are widely utilized in spatial data modeling, notably as priors in generalized mixed effects models (diggle1998model), in spatially varying regression coefficients (gelfand2003spatial), generalized multilevel models (prates2024fast), and in addressing spatial misalignment issues (banerjee2002prediction). We propose the Hausdorff-Gaussian process (HGP), a GP with a correlation function defined on the Hausdorff distance between spatial units. This allows us to directly incorporate shape and size information into the analysis. A crucial aspect of our work is to assess the competitiveness of the HGP, particularly when compared to specialized models designed for areal and mixed spatial data. We investigate whether the HGP, using a powered exponential correlation, can effectively model areal and mixed spatial data, thereby offering a novel and versatile tool for analyzing spatial data with mixed supports.
@@
This paper is structured as follows. Section 2 introduces the Hausdorff distance, discussing its properties, and then delineates the HGP, highlighting its significant applications. Then, in Section 3, we set up spatial generalized linear mixed models (GLMM) with HGP components and detail their inferences in the Bayesian framework. A simulation study to compare the proposed model with specific models for areal and fused data is presented in Section 4. Section 5 demonstrates the practical utility of our model through two distinct applications, one on areal data and the other on data fusion. Section 6 concludes with a discussion. For completeness, we introduce models designed specifically for areal and fused data in the Appendix. The Supplementary Material and the code to reproduce the analyses are available in the following GitHub repository: https://github.com/lcgodoy/jabe-d-25-00205.
@@ 2.1 Hausdorff Distance
The Hausdorff distance is a metric measuring the distance between two sets. Specifically, it is the greatest distance from a point in one set to the closest point in the other set. Consider a metric space $(D,d)$, where $D$ is a spatial region of interest and $d(x,y)$ defines a distance between any two elements $x,y\in D$. For a single point $x$ and a set $A$, define $d(x,A)=\inf_{a\in A}d(x,a)$ as the closest distance from $x$ to any element in $A$. The directed Hausdorff distance from a set $A$ to a set $B$ is $\vec{h}(A,B)=\sup_{a\in A}d(a,B)$. This measure is not necessarily symmetric. The symmetric Hausdorff distance is defined as the greater of the two directed Hausdorff distances:
@@
Note that if $A$ and $B$ are both singletons, then $h(A,B)=d(A,B)$. If they are compact, there exist $a\in A$ and $b\in B$ such that $h(A,B)=d(a,b)$ (arutyunov2016some).
In spatial statistics, sample units (e.g., points, polygons) are repsented as non-empty, closed, and bounded subsets of a study region $D\subset\mathbb{R}^{k}$. In Euclidean domains, the Heine-Borel theorem asserts that closed and bounded sets are compact (abbott2015understanding, pg. 98). The Hausdorff distance is a suitable metric for the collection of these non-empty compact subsets, hereafter denoted $\mathcal{B}(D)$ (molchanov2005theory, Appendix C). By the properties of a metric, $h(A,B)=0$ if and only if $A$ and $B$ are identical. The metric space induced by the tuple $(\mathcal{B}(D),h)$ inherits properties like compactness and completeness from $(D,d)$. Further details regarding the topology induced by the Hausdorff distance can be found in Appendix C in molchanov2005theory. Lastly, in the Supplementary Material, Section S.1, we provide a toy example comparing the Hausdorff distance to the distance between centroids and, in addition, between borders.
@@ 2.2 Hausdorff–Gaussian Process
HGP is a natural extension of GP for modeling spatial data defined on the Hausdorff distance. A GP is a stochastic process over an index set such that all its finite-dimensional marginal distributions are multivariate Normal distributions. Let $D$ be a spatial region of interest. An HGP index set is taken to be the class of non-empty compact subsets of $D$, denoted $\mathcal{B}(D)$. This class encompasses both singletons and polygons, which can be represented as closed and bounded sets. To define an isotropic HGP $Z(\cdot)$ over $\mathcal{B}(D)$, we specify a marginal mean function $m(\cdot)$, a positive-valued marginal standard deviation (SD) function $v(\cdot)$, and a valid isotropic correlation function $r(h)$ for any two sets apart in Hausdorff distance $h$. With these functions, an HGP is a stochastic process $Z(\cdot)$ whose marginal distribution at $\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}\in\mathcal{B}(D)$ is multivariate Normal with mean vector and covariance matrix, respectively,
@@
for $i,j=1,\ldots,n$, where $h(\mathbf{s}_{i},\mathbf{s}_{j})$ is the Hausdorff distance between $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$. This process allows for representations of diverse spatial data, from areal to mixed, making it suitable for a wide range of applications such as disease mapping and data fusion.
Denote the HGP process defined in Equation (1) by $\textrm{HGP}\{m(\cdot),v(\cdot),r(\cdot)\}$. A GP’s mean function has little impact on the process’ local behavior (stein1999interpolation, pg. 12). They capture large-scale variations, and, hereafter $m(\cdot)$ is either held constant or defined through covariates. The SD function $v(\cdot)$ accommodates both homoscedastic and hereroscedastic scenarios. For instance, with a log link, a natural model is
$\log v(\mathbf{s})=\alpha_{0}+\alpha_{1}w(\mathbf{s}),$ (2)
where $w(\mathbf{s})$ is a covariate available for any $s\in\mathcal{B}(D)$. Equation (2) allows a distinction between point-referenced and areal data, a need in data fusion applications (wang2021combining). Specifically, to enable different variances between these data types, we can define $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$, where $\mathcal{A}(\cdot)$ is a function that returns the area of a spatial object. Certainly, care is necessary in specifying $v(\mathbf{s})$ to ensure a valid GP (palacios2006non).
@@ 2.3 Correlation Function
The correlation function $r(\cdot)$ needs to be carefully chosen to ensure the validity of the HGP. In particular, we focus on isotropic correlation functions. Thus, unless explicitly mentioned, all subsequent references to correlation functions will assume they are isotropic. A valid correlation function must be positive definite and meet two key criteria: $r(0)=1$ and $\lim_{h\to\infty}r(h)=0$ (cressie1993statistics, Ch. 2.3.2), where $h$ is the Hausdorff distance between a pair $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{B}(D)$. The latter ensures that spatial correlation diminishes with increasing distance. It is noteworthy that while most geostatistical correlation functions are constructed for Euclidean spaces (cressie1993statistics, Ch. 2.5.1), their positive definiteness may not extend to non-Euclidean spaces (curriero2006use).
@@
where $\nu$ controls the smoothness and $\varphi$ governs the range of the spatial dependence. To enhance interpretability, we reparametrize this function with $\rho={\log(10)}^{1/\nu}\varphi$. Consequently, $\rho$ is the distance at which the spatial correlation between a pair of spatial units drops to $0.10$, commonly referred to as the practical range. The PEC function is appealing because, for $\nu\in(0,1]$, it is positive definite as long as applied to a conditionally negative definite (CND) distance function (schoenberg1938metric). However, to the best of our knowledge, no theoretical results establish when the Hausdorff distance is CND, nor are there results concerning positive definite functions of this distance. Thus, we recommend empirically assessing the positive definiteness of the PEC function before analyzing the data.
To this end, we empirically assess the positive definiteness of the PEC function for $\nu\in(0,1]$. We first calculate Hausdorff distance matrices for the areal and fused data to be analyzed in Sections 5.1 and 5.2, respectively. For comparison purposes, we re-scale these distance matrices so that all their elements are less than or equal to 1. We then evaluate the PEC function for these matrices using a grid of values for $\nu\in(0,1]$ and $\rho\in(0,1.5]$, where the upper bound of $\rho$ represents 1.5 times the largest distance observed in the respective matrix. This process generates a set of corresponding correlation matrices. Finally, we compute the smallest eigenvalue for each of these correlation matrices. A negative eigenvalue indicates the corresponding PEC function is not positive definite for that combination of parameters. However, even valid covariance functions can be ill-conditioned or nearly rank-deficient at dense locations or when there exists a strong spatial dependence (banerjee2013efficient).
@@
Figure 1 displays the smallest eigenvalues ($\lambda_{1}$) of the PEC-induced correlation matrices for the fused and areal datasets from Sections 5.1 and 5.2, respectively. The fused data exhibits positive eigenvalues for all $(\nu,\rho)$ pairs. In contrast, the areal data reveals some small negative eigenvalues (range 0.0099-0.0099 to 0.00006-0.00006) under strong spatial dependence and at $\nu\approx 1$, the border of the parametric space. Based on these empirical results, it is safe to use the PEC function with $\nu\in(0,1]$ for fusion data, but, for the areal application, it may be advisable to stick with a $\nu$ value less than 1. We further investigate the PD property of the PEC function across different datasets and domains. Similiar conclusion are obtained and these results are presented in Section S.2 of the Supplementary Material.
@@ 3.1 Model
Generalized linear models for spatial data can benefit from replacing the GP with the HGP as spatial random effects. In this context, assume $\mathcal{S}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}$ is a set of spatial units such that $\mathbf{s}_{i}\in\mathcal{B}(D)$, where $D\subset\rm I\!R^{2}$. Consider a response variable $Y(\mathbf{s}_{i})$ and a $p$-dimensional vector of covariates $\mathbf{X}_{i}$ associated to the $i$-th spatial unit. Let $\mathbf{Z}={\big[Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n})\big]}^{\top}$ be an $n$-dimensional vector of latent spatial random effects. Then, the spatial GLMM is specified as
$Y(\mathbf{s}_{i})\mid\mathbf{X}_{i},Z(\mathbf{s}_{i})\overset{{\rm ind}}{\sim}f(\cdot\mid\mu_{i},\bm{\gamma}),$
@@
where $g$ is a link function, $\bm{\beta}$ is a $p$-dimensional regression coefficient vector, and $f(\cdot\mid\mu_{i},\bm{\gamma})$ is a probability density or mass function of a distribution with expectation $\mu_{i}$ and an additional parameter vector $\bm{\gamma}$. Thus, conditional on $Z(\mathbf{s}_{i})$ and $\mathbf{X}_{i}$, we assume that the responses $Y(\mathbf{s}_{i})$’s are mutually independent.
The spatial dependence among the responses $\mathbf{Y}={\big[Y(\mathbf{s}_{1}),\ldots,Y(\mathbf{s}_{n})\big]}^{\top}$ is introduced by a zero-mean HGP $\mathbf{Z}\sim\textrm{HGP}\big\{0,v(\cdot;\bm{\sigma}),r(\cdot;\bm{\delta})\big\}$, where $\bm{\sigma}$ and $\bm{\delta}=\{\rho,\nu\}$, are the SD and PEC functions’ parameters, respectively. To summarize, the parameters of the spatial GLMM are $\bm{\theta}={\{\bm{\beta}^{\top},\bm{\sigma}^{\top},\bm{\delta}^{\top},\bm{\gamma}^{\top}\}}^{\top}$.